# Generalizable LLM Learning of Graph Synthetic Data with Reinforcement Learning

### v2

@@ Abstract
Previous research has sought to enhance the graph reasoning capabilities of LLMs by supervised fine-tuning on synthetic graph data. While these led to specialized LLMs better at solving graph algorithm problems, we don’t need LLMs for shortest path: we need generalization from synthetic graph data to real-world tasks with implicit graph structures. In this work, we propose to unlock generalizable learning of graph synthetic data with reinforcement learning. We first design solution-based and process-based rewards for synthetic graph problems: instead of rigid memorizing response patterns in direct fine-tuning, we posit that RL would help LLMs grasp the essentials underlying graph reasoning and alleviate overfitting. We employ RL algorithms such as GRPO and DPO, aligning both off-the-shelf LLMs and LLMs fine-tuned on synthetic graph data. We then compare them against existing settings on both in-domain synthetic tasks and out-of-domain real-world tasks with implicit graph structures such as multi-hop QA, structured planning, and more. Extensive experiments demonstrate that our RL recipe leads to statistically significant improvement on 5 datasets, with an average gain of 12.9% over baseline settings. Further analysis reveals that process-based rewards consistently outperform solution-based rewards, mixing synthetic and real-world task data yields potential gains, while compositionality and explainable intermediate steps remains a critical challenge even after RL.  [^1] Experimental code and results are publicly available at https://anonymous.4open.science/r/Graph_RL-BF08/readme.md
@@ 1 Introduction
Going beyond language processing tasks, large language models (LLMs) are increasingly adopted for tasks with implicit graphical structures, such as commonsense reasoning [63, 62], multi-hop QA [32, 17, 20], and planning [80, 56]. Existing research focuses on evaluating LLMs on the underlying graph problems such as connectivity and shortest path, revealing that LLMs do have preliminary graph reasoning capabilities [84] while suffering from limitations such as robustness and hallucination [84, 100, 25]. Recent works seek to further improve LLM graph reasoning, mostly through supervised fine-tuning (SFT) on graph synthetic data [30, 57]: these techniques demonstrate substantial improvement in LLMs’ ability to tackle graph algorithm problems.
@@
However, we don’t need LLMs to solve synthetic graph problems such as shortest path and topological sort: we already have algorithms (e.g., Dijkstra and Ford-Fulkerson) that are 100% accurate and much more efficient than calling an LLM. The goal of leveraging synthetic graph data should thus be learning from the structured reasoning data and generalizing from synthetic graph problems to real-world tasks with implicit structures. Unfortunately, existing SFT recipes are not improving real-world performance [100], even leading to a 12.5% drop for tasks such as Proscript [63].
@@
To this end, we propose to unlock the power of synthetic graph data with reinforcement learning (RL). Instead of directly memorizing reasoning chains of synthetic problems through SFT, we posit that RL would help LLMs learn the underlying objectives of graph reasoning and mitigate excessive over-fitting to synthetic problems. We design two types of rewards for synthetic graph problems: solution-based, where only the final answer is considered; process-based, where varying weights are given to the soundness of intermediate steps and the final answer. We then align LLMs, both off-the-shelf and fine-tuned on synthetic graph problems, with RL algorithms (e.g., GRPO [67] and DPO [60]) using these reward models. We then compare these RL models against existing settings (off-the-shelf models, SFT, etc.) on 1) in-domain synthetic tasks such as connectivity and shortest path; and more importantly, 2) out-of-domain real-world graph problems such as multi-hop QA [20, 17] and structured commonsense reasoning [62, 63].
Extensive experiments with two LLMs and 8 task settings demonstrate that our proposed RL recipes outperform baselines across tasks and models on both synthetic and real-world graph problems. For synthetic problems, RL models improve by 25% on average for the connectivity task. More importantly, for real-world graph problems, our recipe provides statistically significant improvements on 5 task settings, with an average improvement of 13%. Our experiments also offer insights into the RL recipe: the on-policy algorithm GRPO outperforms the off-policy DPO by 5% on average; process-based rewards offer finer-grained signals for models to learn from when using GRPO, outperforming solution-based by 24% on average on synthetic tasks, while there is no significant edge of process-based synthetic rewards on real-world tasks. Further analysis reveals that mixing a fraction of real-world task data yields potential gains, while the main bottlenecks for RL models are twofold: the compositionality gap from correct single-step results to a correct multi-step solution, and an explainable and hallucination-free intermediate single-step to multi-step reasoning path.
@@ 2 Methdology
For synthetic graph datasets, we follow Wang et al. [84] and Zhang et al. [100] to choose connectivity and shortest path problems. We choose these two tasks since reasoning behind connectivity and shortest path can represent a range of real-world graph problems, including comparing if two ideas are similar or not (if those two ideas are connected) [62], solving an implicit knowledge graph question (where the model will try to reason through the shortest path between two entities) [17]. On the other hand, it is easy and straightforward to verify the correctness of model responses on these two tasks with a Python program.
@@
Our approach follows a three-stage pipeline leveraging synthetic graph data for training and evaluation. We begin with an off-the-shelf model (base model in the following sections) and optionally perform supervised fine-tuning (SFT) on a set of synthetic graph problems. This SFT stage provides the model with exemplars of graph reasoning (including structured reasoning steps and final solutions) in a supervised manner, priming it with domain-specific patterns before reinforcement learning. After SFT (or directly from the base model if SFT is skipped), we further tune the model with reinforcement learning (RL) on synthetic graph tasks. Finally, we evaluate the resulting model on both held-out synthetic problems and real-world graph problems to assess reasoning generalization. This end-to-end pipeline – from base model, to optional synthetic SFT, to RL optimization – is designed to better understand the model’s capabilities of learning and generalizing reasoning on synthetic data to real-world problems.
@@
  • Process-based Reward: We extract the reasoning process and the solution to assign a proper reward to each response, which are enclosed by “<response>...</response>” and “<answer>...</answer>”. For the reasoning processes enclosed by “<response>...</response>”, we reconstruct the underlying graph using NetworkX [27] and evaluate if all the reasoning processes are correct without hallucination or incorrect statements. For the solution, we use the same method as the Solution-based reward to evaluate its correctness.
@@ 3.1 Datasets and Evaluation
We directly adopt the dataset from NLGift [100] and build upon it. For each synthetic task, for each split of train and test sets, we take 500 questions and combine the two synthetic tasks together to construct our synthetic dataset. For the input, we choose to represent the graph using natural language by iterating all nodes first and then for each node iterating through all of the edges (and weights, if applicable). In the instruction, we also specify the desired format for output. For output, we build the ground truth response based on the reward design in Section 2. For evaluation, we use rule-based methods to extract the solution to evaluate the accuracy of each task.
@@
  • Multi-hop QA: Multi-hop QA involves answering questions that need multi-step reasoning on a certain set of knowledge and information, which is strongly related to connectivity (to decide whether two concepts are related) and shortest path (to find the shortest reasoning path to connect two concepts). We adopt StrategyQA [20], Knowledge Crosswords [17], and WikiDYK [101] for the multi-hop QA task.
  • Structured Commonsense Reasoning: Structured commonsense reasoning incooperates commensense into questions that require structured reasoning to solve, which is also related to synthetic tasks like connectivity (to decide if two daily tasks have direct relationship) or shortest path (to decide if two ideas are supporting or countering each other). We adopt ExplaGraphs [62] and Proscript [63] for the structured commonsense reasoning task.
  • Action Planning: Action planning requires following a set of rules and figuring out the correct steps without breaking any of the rules, which is also strongly related to connectivity (to decide if a certain action steps can reach the final goal state) and shortest path (to generate a optimal plan which connects two states). We adopt two versions of Blocksworld [80] (Planning and Verification) for the action planning task.
@@ 3.2 Implementation
We conduct experiments using two base LLMs to verify generality: Qwen2.5-7B-Instruct [93] and Llama-3.1-8B-Instruct [22]. These models are chosen to represent different model families and to ensure our findings are not specific to a single LLM. We chose two reinforcement learning methods for our experiment, one in on-policy RL and one in off-policy RL. For on-policy RL, we use Group Relative Policy Optimization (GRPO) [67], an on-policy reinforcement learning algorithm that forgoes a value critic by comparing groups of model outputs to estimate advantages. For off-policy RL, we apply Direct Preference Optimization (DPO) [60], a reward-alignment technique that fine-tunes the policy on a static dataset of synthetic data generated with the base or tuned model with a high temperature of 0.9. We use a black-box LLM, Gemini 2.0 Flash [58], to summarize the response if needed.
@@
We implement the SFT and DPO training using the HuggingFace TRL library [81], which provides high-level APIs for transformer fine-tuning using SFT and DPO. For GRPO, we utilize the open-source VeRL toolkit [69] to efficiently train the model with distributed rollouts and policy updates. SFT is run for 3 epochs using learning rate of 1e-5, while DPO and GRPO are run for 8 epochs on their respective training data using lr of 1e-6 and 5e-7. We use standard optimization settings and apply the AdamW optimizer in all stages (other hyperparameters are kept consistent across SFT and RL to isolate the effect of training strategy). Training and evaluation are performed on a server with 16 NVIDIA A100 GPUs (40 GB memory each), which allows us to fully fine-tune the 7B/8B models in a reasonable time frame. We use proportions z-test to test statistical significance for all the experiments.
@@ 4 Results
We present main results of our paper in this section. We first evaluate on the held-out set of synthetic problems, and then select the models with good performance on synthetic tasks to evaluate on real-world datasets. While most settings of RL methods achieved significant improvements on synthetic tasks, on real-world implicit graph reasoning tasks there are statistically significant improvements on 5 of the 8 task settings.
@@ 4.1 Synthetic Tasks
As we adopt two reward designs and two RL training methods with an optional SFT stage, for a single model we have eight experiment settings in total. Detailed performance results are shown in Table 1. In general, half of the settings of both models achieved significant improvements (p-val < 0.01), and 3 of the 16 settings achieved similar performance compared to SFT models. Compared with their original base models, training with SFT brings a performance increase of 156% on average, while directly training the model using RL methods provides a performance increase of 120% on average, with GRPO performance better than average and SFT, providing a 222% performance increase. Compared with their SFT models, training additionally with RL brings another 9% performance increase. This indicates that our RL recipes advances synthetic graph problems compared to off-the-shelf models and existing SFT approaches.
@@
RL Method GRPO DPO GRPO DPO
@@
When comparing different RL methods, we see that GRPO achieved better performance across two models with 8 settings, with 7 out of 8 settings achieving significant improvements compared to SFT models, while DPO only achieved significant improvement in 1 out of 8 settings, which is both SFT trained first. The average performance improvement of DPO is lower than those of GRPO, with or without the optional SFT stage. Additionally, GRPO can balance the performance of connectivity and shortest path, resulting in a smaller gap or even better performance on harder shortest path tasks.
@@
SFT stage acts as different roles for different RL methods. For on-policy method GRPO, it may decrease the robustness of the trained model as 2 out 4 settings (SFT + GRPO) cannot achieve comparable results compared to those GRPO trained models trained without SFT stage, leading to limitation of the model’s performance if SFT stage is enforced before GRPO. However, for off-policy method DPO, SFT can somehow bring advantage to the model, with 2 out of 4 settings (SFT + DPO) achieved better results compared to only those only trained using SFT, while those models directly trained using DPO cannot achieve better results than SFT models.
@@ 4.2 Real-World Tasks
Real-world evaluations reveal a helpful but mixed picture: Out of the 8 dataset settings, 3 of them exhibit consistent and statistically significant improvements, and 5 where at least one RL setting is significant. While RL on average achieve positive gains of 13.6%, it still fails to improve on certain settings or tasks. The impact of RL is task-dependent, with benefits most pronounced in problems closer to synthetic tasks such as planning and verification tasks, and some binary answer tasks, but limited improvements or even detrimental effects are observed in certain multi-hop and commonsense reasoning benchmarks. Results are shown in Table 2.
On average, comparing the zero-shot performance, directly synthetic RL using GRPO improved real-world performance by 5.2%, while introducing synthetic SFT before synthetic RL using GRPO improved real-world performance by 20.3%. Compared with synthetic SFT performance, additional training using GRPO and DPO increases performance by 9.5% and 4.0% respectively.
@@
Table 2: Performance of synthetic-only trained models on real-world tasks. Some datasets use abbreviated names: K-C stands for Knowledge Crosswords; WikiDYK-R stands for Reliability setting, while WikiDYK-F stands for factual setting; BW stands for Blocksworld, and BW-P and BW-V stands for planning and verification respectively. -P and -S are reward function choice. All results except Proscript are shown in accuracy, and Proscript result shows percentage of satisfied constraints. Significant performance increases (p-val < 0.01) compared to zero-shot performance are marked bold, and significant performance decreases (p-val < 0.01) are marked with grey. Avg. Increase denotes the average performance increase compared to Zero-Shot. In general, SFT and RL have mixed results on different datasets, while on five of the eight tasks there is at least one RL setting that achieved statistically significant improvement.
@@
A notable observation is that the choice of reward design (process-based vs. solution-based rewards) does not produce a consistent winner across tasks. In some evaluations the process-based reward led to better outcomes, while in others the solution-based reward (which only evaluates the final answer) proved to be more effective. On average, compared with zero-shot performance, tuning with GRPO with process-based or solution-based reward leads to 13% and 8% performance increase when skipping the synthetic SFT stage, while after synthetic SFT the performance increases are 19% and 24% respectively. This inconsistency suggests that the optimal RL reward for synthetic graph problems may be task-dependent.
@@
The impact of purely synthetic SFT is mixed across different datasets. Some tasks derive significant benefit from SFT alone, whereas others show little performance increase or even worse performance. For example, the Blocksworld tasks and WikiDYK factual QA saw immediate boosts from SFT with an average of 56%, which on average RL then amplified performance by 10%. In contrast, other benchmarks did not respond as positively. In tasks like StrategyQA and Knowledge Crosswords, simply applying SFT harmed accuracy, leading to an average of 28% performance decrease comparing to zero-shot performance, which further limits the performance increase using synthetic RL.
@@
Results also suggest that generalization from synthetic to real-world via RL is most successful when the target task is structurally similar to the synthetic tasks used during RL training. In particular, Blocksworld planning task benefited greatly from RL fine-tuning. After training on synthetic planning-like scenarios, the models were far better at solving the Blocksworld planning challenge than zero-shot performance. For instance, an RL-tuned Llama-3.1-8B-Instruct model achieved a dramatically higher planning success rate, increasing 129.7% compared to base or supervised-only counterparts, and an additional 83.2% performance increase when synthetic SFT-trained first.
@@
Finally, we note there are differences between the two model families in how well they generalize with RL. Llama-3.1-8B-Instruct sometimes leverages RL fine-tuning more effectively on certain tasks; however, it also comes with a high variance of decreasing certain tasks’ performance. For instance, compared to a performance increase standard error of 0.1 for Qwen2.5-7B-Instruct trained with GRPO, Llama-3.1-8B-Instruct has a standard error of more than 0.46.
@@
Overall, both RL-trained LLMs show mixed yet encouraging results on real-world tasks with clear successes in certain scenarios, revealing the task-specific nature of RL successes and generalization from synthetic to real-world problems.

### v3

# Generalizable LLM Learning of Graph Synthetic Data with Post-training Alignment

@@ Abstract
Previous research has sought to enhance the graph reasoning capabilities of LLMs by supervised fine-tuning on synthetic graph data. While these led to specialized LLMs better at solving graph algorithm problems, we don’t need LLMs for shortest path: we need generalization from synthetic graph data to real-world tasks with implicit graph structures. In this work, we propose to unlock generalizable learning of graph with post-training alignment with synthetic data. We first design solution-based and process-based rewards for synthetic graph problems: instead of rigid memorizing response patterns in direct fine-tuning, we posit that post-training alignment would help LLMs grasp the essentials underlying graph reasoning and alleviate overfitting on synthetic data. We employ post-training alignment algorithms such as GRPO and DPO, aligning both off-the-shelf LLMs and LLMs fine-tuned on synthetic graph data. We then compare them against existing settings on both in-domain synthetic tasks and out-of-domain real-world tasks with implicit graph structures such as multi-hop QA, structured planning, and more. Extensive experiments demonstrate that our post-training alignment recipe leads to statistically significant improvement on 5 datasets, with an average gain of 12.9% over baseline settings. Further analysis reveals that process-based rewards consistently outperform solution-based rewards on synthetic data but not on real-world tasks, and compositionality and explainable intermediate steps remains a critical challenge even after post-training alignment.  [^1] Experimental code and results are publicly available at https://anonymous.4open.science/r/Graph_RL-BF08/readme.md
Generalizable LLM Learning of Graph Synthetic Data
with Post-training Alignment
Yizhuo Zhangthanks: equal contribution  1  Heng Wang11footnotemark: 1  2  Shangbin Feng11footnotemark: 1  1 Zhaoxuan Tan3 Xinyun Liu4 Yulia Tsvetkov1 1University of Washington   2Xi’an Jiaotong University   3University of Notre Dame   4Google mattyz@uw.edu  wh2213210554@stu.xjtu.edu.cn  shangbin@cs.washington.edu
@@ 1 Introduction
Going beyond language processing tasks, large language models (LLMs) are increasingly adopted for tasks with implicit graphical structures, such as commonsense reasoning (Sakaguchi et al., 2021; Saha et al., 2021), multi-hop QA (Ho et al., 2020; Ding et al., 2024; Geva et al., 2021), and planning (Valmeekam et al., 2023; Padmakumar et al., 2022). Existing research focuses on evaluating LLMs on the underlying graph problems such as connectivity and shortest path, revealing that LLMs do have preliminary graph reasoning capabilities (Wang et al., 2023) while suffering from limitations such as robustness and hallucination (Wang et al., 2023; Zhang et al., 2024b; Guo et al., 2023). Recent works seek to further improve LLM graph reasoning, mostly through supervised fine-tuning (SFT) on graph synthetic data (He et al., 2024b; Perozzi et al., 2024): these techniques demonstrate substantial improvement in LLMs’ ability to tackle graph algorithm problems.
@@
However, we don’t need LLMs to solve synthetic graph problems such as shortest path and topological sort: we already have algorithms (e.g., Dijkstra and Ford-Fulkerson) that are 100% accurate and much more efficient than calling an LLM. The goal of leveraging synthetic graph data should thus be learning from the structured reasoning data and generalizing from synthetic graph problems to real-world tasks with implicit structures. Unfortunately, existing SFT recipes are not improving real-world performance (Zhang et al., 2024b), even leading to a 12.5% drop for tasks such as Proscript (Sakaguchi et al., 2021).
@@
To this end, we propose to unlock the power of synthetic graph data with post-training alignment. Instead of directly memorizing reasoning chains of synthetic problems through SFT, we posit that alignment would help LLMs learn the underlying objectives of graph reasoning and mitigate excessive over-fitting to synthetic problems. We design two types of rewards for synthetic graph problems: solution-based, where only the final answer is considered; process-based, where varying weights are given to the soundness of intermediate steps and the final answer. We then align LLMs, both off-the-shelf and fine-tuned on synthetic graph problems, with alignment algorithms (e.g., GRPO (Shao et al., 2024) and DPO (Rafailov et al., 2023)) using these reward models. We then compare these alignment models against existing settings (off-the-shelf models, SFT, etc.) on 1) in-domain synthetic tasks such as connectivity and shortest path; and more importantly, 2) out-of-domain real-world graph problems such as multi-hop QA (Geva et al., 2021; Ding et al., 2024) and structured commonsense reasoning Saha et al. (2021); Sakaguchi et al. (2021).
@@
Extensive experiments with two LLMs and 8 task settings demonstrate that our proposed alignment recipes can sometimes outperform baselines across tasks and models on both synthetic and real-world graph problems, with important findings on reasoning on gaps between single-step and multi-step reasoning. For synthetic problems, post-training aligned models improve by 25% on average for the connectivity task. More importantly, for real-world graph problems, our recipe provides statistically significant improvements on 5 task settings, with an average improvement of 13%. Our experiments also offer insights into the alignment recipe: the on-policy RL algorithm GRPO outperforms DPO by 5% on average; process-based rewards offer finer-grained signals for models to learn from when using GRPO on synthetic tasks, outperforming solution-based by 24% on average, while there is no significant edge of process-based synthetic rewards on real-world tasks. Further analysis reveals that the main bottlenecks for models after alignment are twofold: the compositionality gap from correct single-step results to a correct multi-step solution, and an explainable and hallucination-free intermediate single-step to multi-step reasoning path.
@@ 2 Methdology
For synthetic graph datasets, we follow Wang et al. (2023) and Zhang et al. (2024b) to choose connectivity and shortest path problems. We choose these two tasks since reasoning behind connectivity and shortest path can represent a range of real-world graph problems, including comparing if two ideas are similar or not (if those two ideas are connected) (Saha et al., 2021), solving an implicit knowledge graph question (where the model will try to reason through the shortest path between two entities) (Ding et al., 2024). On the other hand, it is easy and straightforward to verify the correctness of model responses on these two tasks with a Python program.
@@
Our approach follows a three-stage pipeline leveraging synthetic graph data for training and evaluation. We begin with an off-the-shelf model (base model in the following sections) and optionally perform supervised fine-tuning (SFT) on a set of synthetic graph problems. This SFT stage provides the model with exemplars of graph reasoning (including structured reasoning steps and final solutions) in a supervised manner, priming it with domain-specific patterns before reinforcement learning. After SFT (or directly from the base model if SFT is skipped), we further apply post-training alignment methods on synthetic graph tasks. Finally, we evaluate the resulting model on both held-out synthetic problems and real-world graph problems to assess reasoning generalization. This end-to-end pipeline – from base model, to optional synthetic SFT, to post-training alignment – is designed to better understand the model’s capabilities of learning and generalizing reasoning on synthetic data to real-world problems.
@@
  • Process-based Reward: We extract the reasoning process and the solution to assign a proper reward to each response, which are enclosed by “<response>...</response>” and “<answer>...</answer>”. For the reasoning processes enclosed by “<response>...</response>”, we reconstruct the underlying graph using NetworkX Hagberg et al. (2008) and evaluate if all the reasoning processes are correct without hallucination or incorrect statements. For the solution, we use the same method as the Solution-based reward to evaluate its correctness.
@@ 3.1 Datasets and Evaluation
We directly adopt the dataset from NLGift (Zhang et al., 2024b) and build upon it. For each synthetic task, for each split of train and test sets, we take 500 questions and combine the two synthetic tasks together to construct our synthetic dataset. For the input, we choose to represent the graph using natural language by iterating all nodes first and then for each node iterating through all of the edges (and weights, if applicable). In the instruction, we also specify the desired format for output. For output, we build the ground truth response based on the reward design in Section 2. For evaluation, we use rule-based methods to extract the solution to evaluate the accuracy of each task.
@@
  • Multi-hop QA: Multi-hop QA involves answering questions that need multi-step reasoning on a certain set of knowledge and information, which is strongly related to connectivity (to decide whether two concepts are related) and shortest path (to find the shortest reasoning path to connect two concepts). We adopt StrategyQA (Geva et al., 2021), Knowledge Crosswords (Ding et al., 2024), and WikiDYK (Zhang et al., 2025) for the multi-hop QA task.
  • Structured Commonsense Reasoning: Structured commonsense reasoning incooperates commensense into questions that require structured reasoning to solve, which is also related to synthetic tasks like connectivity (to decide if two daily tasks have direct relationship) or shortest path (to decide if two ideas are supporting or countering each other). We adopt ExplaGraphs (Saha et al., 2021) and Proscript (Sakaguchi et al., 2021) for the structured commonsense reasoning task.
  • Action Planning: Action planning requires following a set of rules and figuring out the correct steps without breaking any of the rules, which is also strongly related to connectivity (to decide if a certain action steps can reach the final goal state) and shortest path (to generate a optimal plan which connects two states). We adopt two versions of Blocksworld (Valmeekam et al., 2023) (Planning and Verification) for the action planning task.
@@ 3.2 Implementation
We conduct experiments using two base LLMs to verify generality: Qwen2.5-7B-Instruct (Yang et al., 2024) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024). These models are chosen to represent different model families and to ensure our findings are not specific to a single LLM. We chose two post-training methods for our experiment, one in on-policy RL and one in supervised learning. For on-policy RL, we use Group Relative Policy Optimization (GRPO) (Shao et al., 2024), an on-policy reinforcement learning algorithm that forgoes a value critic by comparing groups of model outputs to estimate advantages. For supervised supervised learning, we apply Direct Preference Optimization (DPO) (Rafailov et al., 2023), a reward-alignment technique that fine-tunes the policy on a static dataset of synthetic data generated with the base or tuned model with a high temperature of 0.9. We use a black-box LLM, Gemini 2.0 Flash (Pichai, 2024), to summarize the response if needed.
@@
We implement the SFT and DPO training using the HuggingFace TRL library (von Werra et al., 2020), which provides high-level APIs for transformer fine-tuning using SFT and DPO. For GRPO, we utilize the open-source VeRL toolkit (Sheng et al., 2024) to efficiently train the model with distributed rollouts and policy updates. SFT is run for 3 epochs using learning rate of 1e-5, while DPO and GRPO are run for 8 epochs on their respective training data using lr of 1e-6 and 5e-7. We use standard optimization settings and apply the AdamW optimizer in all stages (other hyperparameters are kept consistent across SFT and RL to isolate the effect of training strategy). Training and evaluation are performed on a server with 16 NVIDIA A100 GPUs (40 GB memory each), which allows us to fully fine-tune the 7B/8B models in a reasonable time frame. We use proportions z-test to test statistical significance for all the experiments.
@@ 4 Results
We present main results of our paper in this section. We first evaluate on the held-out set of synthetic problems, and then select the models with good performance on synthetic tasks to evaluate on real-world datasets. While most settings of alignment methods achieved significant improvements on synthetic tasks, on real-world implicit graph reasoning tasks there are statistically significant improvements on 5 of the 8 task settings.
@@ 4.1 Synthetic Tasks
As we adopt two reward designs and two alignment training methods with an optional SFT stage, for a single model we have eight experiment settings in total. Detailed performance results are shown in Table 1. In general, half of the settings of both models achieved significant improvements (p-val < 0.01), and 3 of the 16 settings achieved similar performance compared to SFT models. Compared with their original base models, training with SFT brings a performance increase of 156% on average, while directly training the model using alignment methods provides a performance increase of 120% on average, with GRPO performance better than average and SFT, providing a 222% performance increase. Compared with their SFT models, training additionally with alignment brings another 9% performance increase. This indicates that our alignment recipes advances synthetic graph problems compared to off-the-shelf models and existing SFT approaches.
@@
Alignment Method GRPO DPO GRPO DPO
@@
When comparing different alignment methods, we see that GRPO achieved better performance across two models with 8 settings, with 7 out of 8 settings achieving significant improvements compared to SFT models, while DPO only achieved significant improvement in 1 out of 8 settings, which is both SFT trained first. The average performance improvement of DPO is lower than those of GRPO, with or without the optional SFT stage. Additionally, GRPO can balance the performance of connectivity and shortest path, resulting in a smaller gap or even better performance on harder shortest path tasks.
@@ 4.2 Real-World Tasks
Real-world evaluations reveal a helpful but mixed picture: Out of the 8 dataset settings, 3 of them exhibit consistent and statistically significant improvements, and 5 where at least one alignment setting is significant. While post-training alignment on average achieve positive gains of 13.6%, it still fails to improve on certain settings or tasks. The impact of post-training alignment is task-dependent, with benefits most pronounced in problems closer to synthetic tasks such as planning and verification tasks, and some binary answer tasks, but limited improvements or even detrimental effects are observed in certain multi-hop and commonsense reasoning benchmarks. Results are shown in Table 2.
On average, comparing the zero-shot performance, directly synthetic alignment using GRPO improved real-world performance by 5.2%, while introducing synthetic SFT before synthetic alightnment using GRPO improved real-world performance by 20.3%. Compared with synthetic SFT performance, additional training using GRPO and DPO increases performance by 9.5% and 4.0% respectively.
@@
Table 2: Performance of synthetic-only trained models on real-world tasks. Some datasets use abbreviated names: K-C stands for Knowledge Crosswords; WikiDYK-R stands for Reliability setting, while WikiDYK-F stands for factual setting; BW stands for Blocksworld, and BW-P and BW-V stands for planning and verification respectively. -P and -S are reward function choice. All results except Proscript are shown in accuracy, and Proscript result shows percentage of satisfied constraints. Significant performance increases (p-val < 0.01) compared to zero-shot performance are marked bold, and significant performance decreases (p-val < 0.01) are marked with grey. Avg. Increase denotes the average performance increase compared to Zero-Shot. In general, SFT and post-training alignment have mixed results on different datasets, while on five of the eight tasks there is at least one alignment setting that achieved statistically significant improvement.
@@
A notable observation is that the choice of reward design (process-based vs. solution-based rewards) does not produce a consistent winner across tasks. In some evaluations the process-based reward led to better outcomes, while in others the solution-based reward (which only evaluates the final answer) proved to be more effective. On average, compared with zero-shot performance, tuning with GRPO with process-based or solution-based reward leads to 13% and 8% performance increase when skipping the synthetic SFT stage, while after synthetic SFT the performance increases are 19% and 24% respectively. This inconsistency suggests that the optimal alignment reward for synthetic graph problems may be task-dependent.
@@
The impact of purely synthetic SFT is mixed across different datasets. Some tasks derive significant benefit from SFT alone, whereas others show little performance increase or even worse performance. For example, the Blocksworld tasks and WikiDYK factual QA saw immediate boosts from SFT with an average of 56%, which on average alignment then amplified performance by 10%. In contrast, other benchmarks did not respond as positively. In tasks like StrategyQA and Knowledge Crosswords, simply applying SFT harmed accuracy, leading to an average of 28% performance decrease comparing to zero-shot performance, which further limits the performance increase using synthetic alignment.
@@
Results also suggest that generalization from synthetic to real-world via alignment is most successful when the target task is structurally similar to the synthetic tasks used during post-training alignment. In particular, Blocksworld planning task benefited greatly from alignment. After training on synthetic planning-like scenarios, the models were far better at solving the Blocksworld planning challenge than zero-shot performance. For instance, an alignment-tuned Llama-3.1-8B-Instruct model achieved a dramatically higher planning success rate, increasing 129.7% compared to base or supervised-only counterparts, and an additional 83.2% performance increase when synthetic SFT-trained first.
@@
Finally, we note there are differences between the two model families in how well they generalize with alignment. Llama-3.1-8B-Instruct sometimes leverages post-training alignment more effectively on certain tasks; however, it also comes with a high variance of decreasing certain tasks’ performance. For instance, compared to a performance increase standard error of 0.1 for Qwen2.5-7B-Instruct trained with GRPO, Llama-3.1-8B-Instruct has a standard error of more than 0.46.
@@
Overall, both post-training aligned LLMs show mixed yet encouraging results on real-world tasks with clear successes in certain scenarios, revealing the task-specific nature of post-training alignment successes and generalization from synthetic to real-world problems.